OpenStudy (anonymous):

32/? = 20/25 what is ?

OpenStudy (math&ing001):

When faced with these kinda problems try the rule of 3 32 -> 20 ? -> 25 your number would be 25*32/20

OpenStudy (unklerhaukus):

a/b = c/d a = cb/d da = cb da/c = b

OpenStudy (texaschic101):

32/x = 20/25 This is a proportion. To solve it, you cross multiply and solve for x. (20)(x) = (32)(25) multiply through and solve for x. You can then check your answer by subbing in what you got for x back into the proportion, cross multiply, and if both sides come out equal, your answer is correct.